A car starts from rest and acquires a velocity of 70 km/h in 20 minutes. Find a) the acceleration; b) the distance traveled in 20 minutes, and c) the distance traveled in the next 20 minutes.



1
Expert's answer
2021-09-20T14:42:14-0400

"70\ km/h=19.4\ m/s"


"20\ min=1200\ s"


a) "v=at\to a=v/t=19.4/1200=0.016\ (m/s^2)"


b) "s=at^2/2=0.016\cdot1200^2/2=11640\ (m)"


c) "s_1=v_0t+at^2/2=19.4\cdot1200+0.016\cdot1200^2/2=34800\ (m)"